OpenStudy (anonymous):

Hi can you help me with the odds on these problems? Could you also do it step by step and tell me which formula or identity you used? Thank you so much

OpenStudy (anonymous):

Here is another one. Oh and this is precalc ilia trigonometry

ganeshie8 (ganeshie8):

start by writing the left hand side as single sin or cos function by using angle sum identitites : \[\begin{align} \sin\theta - \sqrt{3}\cos \theta &= \frac{2}{2}\left[\sin\theta - \sqrt{3}\cos \theta\right]\\~\\ &= 2\left[\sin\theta~\frac{1}{2} -\cos \theta~\frac{\sqrt{3}}{2}\right]\\~\\ &= 2\left[\sin\theta~\cos ? - \cos \theta~\sin ?\right]\\~\\ \end{align}\]

OpenStudy (anonymous):

Okay then what's next?:)